OpenStudy (anonymous):

A carpenter has two pieces of oak trim. One piece of trim is 3 3/4ft long and the other is 4 3/8ft long. Estimate how many feet of oak trim she has in all?

OpenStudy (anonymous):

33/4 ft or, 15/4 ft

OpenStudy (anonymous):

so 15/4

OpenStudy (anonymous):

estimate would be 16

OpenStudy (anonymous):

how come

OpenStudy (anonymous):

idk?

Parth (parthkohli):

@Jyako90 and everyone, it's a mixed fraction, you misunderstood the question.

Parth (parthkohli):

4 + 4 is the estimate. = 8

Parth (parthkohli):

8 is the answer in estimation

OpenStudy (anonymous):

60/8 27/8

Parth (parthkohli):

15/4 + 35/8 = 65/8 = 8 1/8 .. so estimation is 8.

OpenStudy (anonymous):

60/8+27/8

OpenStudy (anonymous):

87/8? is that right

OpenStudy (anonymous):

see he has one piece = \[ {{15}\over{4}}ft\] see he has another piece = \[ {{35}\over{8}}ft\]

OpenStudy (anonymous):

add these and see how much is with him

OpenStudy (anonymous):

i did 87/8 the denominator has to be same to add

OpenStudy (anonymous):

it should be \[{30+35} \over{8}\] = \[{{65} \over{8}}ft\]

OpenStudy (anonymous):

no no sorry it was my mistake i took 3 first but it is \[4{{3}\over{8}}ft ={ {4\times8+3}\over{8}}ft = {{35}\over{8}}ft\]